The vertex of the function
f(x) = x^2 + 6x +7

Answer by user_dude2008(1862)

x=-b/2a=(-6)/(2*1) = -3

f(-3)=(-3)^2+6(-3)+7=9-18+7=-9+7=-2

Answer: Vertex is (-3,-2)